# Regions of suppressed diffusion around supernova remnants?

### v1

@@ 1 Introduction
These questions are not all independent from each other: for instance, the answer to the third question depends very sensibly upon whether the diffusion coefficient is only small around PWNe or also around SNRs, and it also depends on whether the diffusion coefficient is small everywhere in the disc or only in cocoons of limited size around sources. Moreover, the possibility for CRs to excite streaming instability and lead to reduced diffusivity depends on whether the particles are protons from a SNR or rather a quasi-neutral electron-positron beam from PWNe, or even a spatially charge-separated beam of very high energy electrons or positrons from PWNe Olmi & Bucciantini (2019); Olmi2024.
@@
In order to start addressing these issues, we consider two approaches to CR diffusion. The first is a simple phenomenological one already used by several authors before (for instance Aharonian et al. (2008); Nava & Gabici (2013); Bao2019; Bao2021), in which diffusion is assumed to occur in one or three spatial dimensions, at a level that can be calibrated to fit the data. In other words the normalisation and energy dependence of the diffusion coefficient are parameters of the problem. This approach is appropriate to describe transport in a disc where the coherence length is small or there is an enhanced level of turbulence, so that spatial diffusion is slow. The approach may also be used to check the possibility that some set of data may hint at small diffusion coefficient, independent of the origin of the suppressed diffusion.
@@ 2 Differences and Analogies between PWNe and SNRs as Particle Accelerators
On the other hand, the excitation of streaming instability by the pairs escaping a PWN did not receive equal attention: if the pairs are liberated into the ISM without charge separation, the total current that they carry vanishes and the non resonant branch (Bell, 2004) of the instability cannot be excited. The resonant streaming instability is excited even in a situation of zero net current (Evoli et al., 2018) but the resulting level of self-generated turbulence is insufficient to explain the strong suppression of diffusion coefficient, unless the transverse cross section of the flux tube is assumed to be very small (Mukhopadhyay & Linden, 2022), which is clearly at odds with the quasi-spherical morphology of the emission. Some evidence that electrons and positrons at energies comparable with the potential drop of the pulsar may be injected at different locations, thereby making the net current locally non-zero, was presented by (Olmi & Bucciantini, 2019). The effect of the resulting current for the excitation of the non-resonant streaming instability is currently being investigated Olmi2024. This phenomenon might be related to another non-thermal unexpected occurrence, namely the long and narrow bright X-ray filaments seen around some PWNe (see for instance Hui & Becker (2007)). This latter phenomenon might be somehow related to the production of TeV halos at later stages.
@@ 3.1 HB9
Assuming simple dynamical evolution of the SNR in a homogeneous medium, the maximum energy of the accelerated particles in HB9 can be estimated following Equation (21) in Bell et al. (2013) or Equation (27) in Cardillo et al. (2015). The maximum energy was $\sim 20$ TeV at the beginning of the ST phase and is $\sim 5$ TeV at present. These estimates can be appreciably lower if damping happens to play an important role. This may be the case especially at late times, when the non resonant hybrid instability can no longer be excited, and ion-neutral damping may be in action, making the maximum energy of the accelerated particles appreciably lower than estimated above (see for instance Fig. 2 of ZP2003).
@@ 4.1 Phenomenological diffusion coefficient
The phenomenological approach to the description of CR transport around the SNRs HB9 and W28 is based upon the assumption that the diffusion coefficient around these sources is spatially constant and numerically estimated to be a fraction of the Galactic diffusion coefficient $D_{\rm Gal}$, as estimated from the secondary/primary ratios (see e.g., Ptuskin et al., 2009). The exact numerical value of $D_{\rm Gal}$ is not very important in this context, since it is only used as a reference with respect to which the reduced diffusivity manifests itself. Thus, we assume $D(p)=D_{\rm Gal}/\kappa$, with the suppression factor $\kappa>1$, $p$ the proton momentum, and for the Galactic diffusion coefficient we use $D_{\rm Gal}=3.6\times 10^{28}[{\rm pc}/(1\ {\rm GeV{/c}})]^{1/3}$ cm2 s-1, aware that deviations from this trend are likely, especially at low energies Evoli et al. (2020). This simple approach is adopted here following the same logic as in the past literature on the topic Aharonian et al. (2008); Nava & Gabici (2013); Bao2019; Bao2021, namely to have a quick and approximate assessment of the possible suppression of diffusivity in the near-source regions.
@@ 4.2 Self-generated diffusion
In the region surrounding a source we can naturally expect CR energy density in excess of the mean Galactic value, as well as stronger CR gradients that result from the presence of a source. In these conditions, as previously discussed by (Ptuskin et al., 2008; Malkov2013; D’Angelo et al., 2016; Nava et al., 2016; D’Angelo et al., 2018; Nava et al., 2019; Recchia et al., 2022), the transport of CRs may be dominated by self-generated perturbations, mainly in the form of Alfvén waves moving parallel to the regular Galactic magnetic field in which the source is assumed to be located. Since the gradient has a well defined sign, most Alfvén waves move away from the source. This results in an advective term in the transport equation with an advection velocity that equals the Alfvén speed. The main novelty of the present calculations is the fact that we account for the possible presence of dense MCs located in the regions around specific sources and with properties that may lead to the identification of reduced diffusivity. Moreover we investigate the effect of the MCs in terms of grammage accumulated by CRs while propagating away from the sources.
@@ 5.1 Phenomenological models of CR transport
The issue of whether it is reasonable to have a maximum energy in the $\sim 100$ GeV range for a SNR having the age of HB9 is strongly related to the damping mechanisms that are at work in the acceleration region. For instance, in Fig. 2 of ZP2003 one can read the maximum energy as a function of time in the presence of damping. If only non-linear Landau damping is at work, it is not feasible to decrease the maximum energy to such low values. If ion-neutral damping near the shock dominates, a maximum energy below $\sim 1$ TeV seems somewhat too low for HB9, but it is hard to rule it out, given the uncertainties on the local density and the ionization fraction in the acceleration region. Hence, unfortunately some level of degeneracy between the two scenarios discussed above cannot be excluded.
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 5.2 Models of self-generated transport
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Taken at face value, this finding would be clear evidence that even in the case of an under-luminous SNR such as HB9 there is suppression of diffusivity around the source. Unfortunately, as discussed in Sec. 5.1, the low energy $\gamma$-ray flux reduction suggesting a smaller diffusivity can also be caused by the differential escape of CRs from the upstream region of the shock, if the maximum energy of the accelerated particles for a SNR with the age of HB9 is as low as $\sim 100$ GeV. Although this condition requires a role of ion-neutral damping at the shock that is probably larger than what one would expect ZP2003, the scenario cannot be ruled out given the uncertainties in the environmental parameters.
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 7 Conclusion and discussion
A rough assessment of the possible need for a suppression in the diffusion coefficient around sources can be made by using a phenomenological calculation, in which the diffusion coefficient is parametrized as a fraction of a Galactic-like diffusion coefficient Aharonian et al. (2008); Nava & Gabici (2013); Bao2019; Bao2021. Meanwhile, a non-linear computation that includes self-generation and damping, clarifies whether the conditions are appropriate to expect a suppression of the diffusion coefficient as a result of the excitation of resonant streaming instability.

### v2

@@ 1 Introduction
These questions are not all independent from each other: for instance, the answer to the third question depends very sensibly upon whether the diffusion coefficient is only small around PWNe or also around SNRs, and it also depends on whether the diffusion coefficient is small everywhere in the disc or only in cocoons of limited size around sources. Moreover, the possibility for CRs to excite streaming instability and lead to reduced diffusivity depends on whether the particles are protons from a SNR or rather a quasi-neutral electron-positron beam from PWNe, or even a spatially charge-separated beam of very high energy electrons or positrons from PWNe Olmi & Bucciantini (2019); Olmi et al. (2024).
@@
In order to start addressing these issues, we consider two approaches to CR diffusion. The first is a simple phenomenological one already used by several authors before (for instance Aharonian et al. (2008); Nava & Gabici (2013); Bao et al. (2019); Bao & Chen (2021)), in which diffusion is assumed to occur in one or three spatial dimensions, at a level that can be calibrated to fit the data. In other words the normalisation and energy dependence of the diffusion coefficient are parameters of the problem. This approach is appropriate to describe transport in a disc where the coherence length is small or there is an enhanced level of turbulence, so that spatial diffusion is slow. The approach may also be used to check the possibility that some set of data may hint at small diffusion coefficient, independent of the origin of the suppressed diffusion.
@@ 2 Differences and Analogies between PWNe and SNRs as Particle Accelerators
On the other hand, the excitation of streaming instability by the pairs escaping a PWN did not receive equal attention: if the pairs are liberated into the ISM without charge separation, the total current that they carry vanishes and the non resonant branch (Bell, 2004) of the instability cannot be excited. The resonant streaming instability is excited even in a situation of zero net current (Evoli et al., 2018) but the resulting level of self-generated turbulence is insufficient to explain the strong suppression of diffusion coefficient, unless the transverse cross section of the flux tube is assumed to be very small (Mukhopadhyay & Linden, 2022), which is clearly at odds with the quasi-spherical morphology of the emission. Some evidence that electrons and positrons at energies comparable with the potential drop of the pulsar may be injected at different locations, thereby making the net current locally non-zero, was presented by (Olmi & Bucciantini, 2019). The effect of the resulting current for the excitation of the non-resonant streaming instability is currently being investigated Olmi et al. (2024). This phenomenon might be related to another non-thermal unexpected occurrence, namely the long and narrow bright X-ray filaments seen around some PWNe (see for instance Hui & Becker (2007)). This latter phenomenon might be somehow related to the production of TeV halos at later stages.
@@ 3.1 HB9
Assuming simple dynamical evolution of the SNR in a homogeneous medium, the maximum energy of the accelerated particles in HB9 can be estimated following Equation (21) in Bell et al. (2013) or Equation (27) in Cardillo et al. (2015). The maximum energy was $\sim 20$ TeV at the beginning of the ST phase and is $\sim 5$ TeV at present. These estimates can be appreciably lower if damping happens to play an important role. This may be the case especially at late times, when the non resonant hybrid instability can no longer be excited, and ion-neutral damping may be in action, making the maximum energy of the accelerated particles appreciably lower than estimated above (see for instance Fig. 2 of Ptuskin & Zirakashvili (2003)).
@@ 4.1 Phenomenological diffusion coefficient
The phenomenological approach to the description of CR transport around the SNRs HB9 and W28 is based upon the assumption that the diffusion coefficient around these sources is spatially constant and numerically estimated to be a fraction of the Galactic diffusion coefficient $D_{\rm Gal}$, as estimated from the secondary/primary ratios (see e.g., Ptuskin et al., 2009). The exact numerical value of $D_{\rm Gal}$ is not very important in this context, since it is only used as a reference with respect to which the reduced diffusivity manifests itself. Thus, we assume $D(p)=D_{\rm Gal}/\kappa$, with the suppression factor $\kappa>1$, $p$ the proton momentum, and for the Galactic diffusion coefficient we use $D_{\rm Gal}=3.6\times 10^{28}[{\rm pc}/(1\ {\rm GeV{/c}})]^{1/3}$ cm2 s-1, aware that deviations from this trend are likely, especially at low energies Evoli et al. (2020). This simple approach is adopted here following the same logic as in the past literature on the topic Aharonian et al. (2008); Nava & Gabici (2013); Bao et al. (2019); Bao & Chen (2021), namely to have a quick and approximate assessment of the possible suppression of diffusivity in the near-source regions.
@@ 4.2 Self-generated diffusion
In the region surrounding a source we can naturally expect CR energy density in excess of the mean Galactic value, as well as stronger CR gradients that result from the presence of a source. In these conditions, as previously discussed by (Ptuskin et al., 2008; Malkov et al., 2013; D’Angelo et al., 2016; Nava et al., 2016; D’Angelo et al., 2018; Nava et al., 2019; Recchia et al., 2022), the transport of CRs may be dominated by self-generated perturbations, mainly in the form of Alfvén waves moving parallel to the regular Galactic magnetic field in which the source is assumed to be located. Since the gradient has a well defined sign, most Alfvén waves move away from the source. This results in an advective term in the transport equation with an advection velocity that equals the Alfvén speed. The main novelty of the present calculations is the fact that we account for the possible presence of dense MCs located in the regions around specific sources and with properties that may lead to the identification of reduced diffusivity. Moreover we investigate the effect of the MCs in terms of grammage accumulated by CRs while propagating away from the sources.
@@ 5.1 Phenomenological models of CR transport
The issue of whether it is reasonable to have a maximum energy in the $\sim 100$ GeV range for a SNR having the age of HB9 is strongly related to the damping mechanisms that are at work in the acceleration region. For instance, in Fig. 2 of Ptuskin & Zirakashvili (2003) one can read the maximum energy as a function of time in the presence of damping. If only non-linear Landau damping is at work, it is not feasible to decrease the maximum energy to such low values. If ion-neutral damping near the shock dominates, a maximum energy below $\sim 1$ TeV seems somewhat too low for HB9, but it is hard to rule it out, given the uncertainties on the local density and the ionization fraction in the acceleration region. Hence, unfortunately some level of degeneracy between the two scenarios discussed above cannot be excluded.
@@ 5.2 Models of self-generated transport
Taken at face value, this finding would be clear evidence that even in the case of an under-luminous SNR such as HB9 there is suppression of diffusivity around the source. Unfortunately, as discussed in Sec. 5.1, the low energy $\gamma$-ray flux reduction suggesting a smaller diffusivity can also be caused by the differential escape of CRs from the upstream region of the shock, if the maximum energy of the accelerated particles for a SNR with the age of HB9 is as low as $\sim 100$ GeV. Although this condition requires a role of ion-neutral damping at the shock that is probably larger than what one would expect Ptuskin & Zirakashvili (2003), the scenario cannot be ruled out given the uncertainties in the environmental parameters.
@@ 7 Conclusion and discussion
A rough assessment of the possible need for a suppression in the diffusion coefficient around sources can be made by using a phenomenological calculation, in which the diffusion coefficient is parametrized as a fraction of a Galactic-like diffusion coefficient Aharonian et al. (2008); Nava & Gabici (2013); Bao et al. (2019); Bao & Chen (2021). Meanwhile, a non-linear computation that includes self-generation and damping, clarifies whether the conditions are appropriate to expect a suppression of the diffusion coefficient as a result of the excitation of resonant streaming instability.